Bowman Specialists Inc. (BSI) manufactures specialized equipment for polishing optical lenses. There are two models—one (A–25) principally used for fine eyewear and the other (A–10) for lenses used in binoculars, cameras, and similar equipment.

The following table shows the manufacturing cost of each unit is calculated, using activity-based costing, for these manufacturing cost pools.

Cost Pools Allocation Base Costing Rate
Materials handling Number of parts $ 3.75 per part
Manufacturing supervision Hours of machine time $ 24.50 per hour
Assembly Number of parts $ 5.55 per part
Machine setup Each setup $ 48.60 per setup
Inspection and testing Logged hours $ 65.00 per hour
Packaging Logged hours $ 35.00 per hour

BSI currently sells the A–10 model for $3,250 and the A–25 model for $1,835. Manufacturing costs and activity usage for the two products follow:

A-10 A-25
Direct materials $ 143.76 $ 76.44
Number of parts 131 102
Machine-hours 9.00 6.00
Inspection time 2.00 1.10
Packing time 1.20 0.60
Setups 22 11

Required:

1. Calculate the product cost and product margin for each product. (Round answers to 2 decimal places)

2. A new competitor has entered the market for lens-polishing equipment with a superior product at significantly lower prices, $2,330 for the A–10 model and $1,715 for the A–25 model. To try to compete, BSI has made some radical improvements in the design and manufacturing of its two products. The materials costs and activity usage rates have been decreased significantly, as follows:

A-10 A-25
Direct materials $ 98.65 $ 52.45
Number of parts 130 101
Machine-hours 10.0 4.0
Inspection time 2.0 1.00
Packing time 1.00 0.40
Setups 11 11

2-a. Calculate the total product costs with the new activity usage data. (Round answers to 2 decimal places)

2-b. Can BSI make a positive gross margin with the new costs, assuming that it must meet the price set by the new competitor? Yes or No?

4. What cost management method might be useful to BSI at this time? Target costing, Activity-based costing, Life-cycle costing or Kaizen costing?
Answer #1
1.
A-10 ($) A-25 ($)
Direct Material      143.76         76.44
Material Handling =131*3.75      491.25 =102*3.75      382.50
Manufacturing Supervision =9*24.5      220.50 =6*24.5      147.00
Assembly =131*5.55      727.05 =102*5.55      566.10
Machine Setup =22*48.6 1,069.20 =11*48.6      534.60
Inspection and testing =2*65      130.00 =1.1*65         71.50
Packaging =1.2*35         42.00 =0.6*35         21.00
Total product cost per unit (A) 2,823.76 1,799.14
Selling price per unit ($) (B) 3,250.00 1,835.00
Product Margin per unit (B-A)= C      426.24         35.86
2-a.
A-10 ($) A-25 ($)
Direct Material         98.65         52.45
Material Handling =130*3.75      487.50 =101*3.75      378.75
Manufacturing Supervision =10*24.5      245.00 =4*24.5         98.00
Assembly =130*5.55      721.50 =101*5.55      560.55
Machine Setup =11*48.6      534.60 =11*48.6      534.60
Inspection and testing =2*65      130.00 =1*65         65.00
Packaging =1*35         35.00 =0.4*35         14.00
Total product cost per unit (A) 2,252.25 1,703.35
2-b.
A-10 ($) A-25 ($)
Selling price per unit ($) (A)    2,330.00 1,715.00
Cost per unit (B)    2,252.25 1,703.35
Product Margin per unit (A-B)= C          77.75         11.65
Yes, BSI make positive gross margin with the new cost and price set by new cmpetitor
4.
Kaizen costing as competitor set up new selling price which are less than earlier prices, in this costing system there is continuous imoprovement in cost reduction which will ultimately increase the product margin.